OpenStudy (lovelyharmonics):

Find the first six terms of the sequence. a1 = -3, an = 2 ● an-1

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

what is that big dot?

OpenStudy (anonymous):

\[\large a_1=-3,a_n=2a_{n-1}\]

OpenStudy (lovelyharmonics):

lol idk i just copied and pasted the problem XD i think its a multiplication sign though

OpenStudy (anonymous):

this means the first term is \(-3\) and to get to the next term you multiply by \(2\)

OpenStudy (lovelyharmonics):

.-. so i subtract 3 from a number and then multiply what i get by 2?

OpenStudy (anonymous):

in other words,\[a_1=-3,\\a_2=2\times (-3)=-6\\ a_3=2\times (-6)=-12\] and so on

OpenStudy (anonymous):

no don't subtract, just keep multiplying by \(2\)

OpenStudy (lovelyharmonics):

oh so then its -3, -6, -12, -24, -48, -96?

OpenStudy (anonymous):

yup

OpenStudy (lovelyharmonics):

wait would it be that or this? -6, -12, -24, -48, -96, -192

OpenStudy (anonymous):

hope it is clear from the subscripts \[a_n=2a_{n-1}\] means the next term is \(2\) times the previous one

OpenStudy (anonymous):

says "first six" and the first one is \(-3\)

OpenStudy (lovelyharmonics):

okay but how do i do it backwards such as this? Find an explicit rule for the nth term of the sequence. 2, -8, 32, -128, ...

OpenStudy (anonymous):

looks to me like you are multiplying by \(-4\) each time

OpenStudy (anonymous):

so you could say \[a_1=2,a_n=-4\times a_{n-1}\]

OpenStudy (anonymous):

or use a big ugly dot instead of the \(\times\)

OpenStudy (lovelyharmonics):

XD hey i think its a cute dot

OpenStudy (anonymous):

\[\bullet\] \[\odot\]

OpenStudy (lovelyharmonics):

couldnt you also write that as 2 x (-4)^n-1

OpenStudy (anonymous):

no

OpenStudy (anonymous):

oh, doh, yes!

OpenStudy (lovelyharmonics):

yay c:

OpenStudy (anonymous):

that is the closed form i wrote the recursion

OpenStudy (lovelyharmonics):

.-. no idea what you just said

OpenStudy (anonymous):

recursion is the one where you say what to do to the previous term to get the next term

OpenStudy (lovelyharmonics):

oh that makes sense c: okay thank you c:

OpenStudy (anonymous):

\[a_1=2,a_n=-4\times a_{n-1}\] is a recursion the one you wrote \[a_n=2\times (-4)^{n-1}\] is the "closed form" meaning if you just plug in \(n\)

OpenStudy (anonymous):

yw